A fragment of DNA is amplified from a single molecule of template DNA for 22 cycles.

Theoretically, how many PCR products will be produced?. Assume exponential amplification starts with cycle 1.

please explain and how would the answer be different if only one primer was used?
Answer #1

Number of copies of DNA molecule formed after n cycles of PCR = 2^n

Here, n = 22

Number of DNA copies = 2^22 = 4,194,304

This is exponential increase in the number of copies of DNA molecule. If only one primer is used, than the number of copies of DNA molecule will increase in a linear way and a total of 22 copies of DNA will be formed.

Please give high rating.

​​​​​